# The influence of AGN feedback on star formation in red spiral galaxies

### v1

@@ 4 Discussion
Our findings are consistent with the dual role of AGN feedback in spiral galaxies. Optical emission-line diagnostics reveal that a significant fraction of red spirals exhibit AGN and LINER signatures, consistent with negative feedback processes that suppress star formation and contribute to quenching. This aligns with previous studies (e.g. Schawinski et al., 2007; Cid Fernandes et al., 2011; Povic et al., 2012), which argue that AGNs are instrumental in transitioning galaxies across the green valley toward the red sequence.
@@
In Figure 10, we show the redshift vs. cluster-centric radius () of red and blue galaxies classified as “star-forming,” “AGN,” and “LINER” systems. In the top panel of Figure 10, galaxy classification is based on the standard BPT plot using [O III]/ vs. [N II]/ (see Figure 2). In the middle panel, the [O III]/ vs. [S II]/ emission line ratios indicate activity (see Figure 3). Finally, the bottom panel depicts galaxies classified based on activity indicated by [O III]/ vs. [O I]/ (see Figure 4)). For all three panels, projected cluster-centric distance was normalized to (i.e., ), the radius of a sphere whose density is 200 times the critical density of the universe (cf. Barkhouse et al., 2007).
@@
For a complementary analysis of stellar population age and star formation activity, including Dn(4000) and specific star formation rate metrics, we refer to Paper I which explores the UV and IR properties of our galaxy sample.
@@ 5 Conclusions
A majority of red spiral galaxies display AGN or LINER signatures in optical diagnostics, consistent with feedback-driven suppression of star formation.
@@
A significant fraction of blue spirals shows high X-ray luminosities despite being classified as star-forming or composite (star-forming/AGN) in optical diagrams, indicating hidden or optically obscured AGN activity.

### v2

@@ 4 Discussion
Our findings are consistent with the dual role of AGN feedback in spiral galaxies. Optical emission-line diagnostics reveal that a significant fraction (71% to 97%) of red spirals exhibit AGN and LINER signatures, consistent with negative feedback processes that suppress star formation and contribute to quenching. This aligns with previous studies (e.g. Schawinski et al., 2007; Cid Fernandes et al., 2011; Povic et al., 2012), which argue that AGNs are instrumental in transitioning galaxies across the green valley toward the red sequence.
@@
In Figure 10, we show the redshift vs. cluster-centric radius () of red and blue galaxies classified as “star-forming,” “AGN,” and “LINER” systems. In the top panel of Figure 10, galaxy classification is based on the standard BPT plot using [O III]/ vs. [N II]/ (see Figure 2). In the middle panel, the [O III]/ vs. [S II]/ emission line ratios indicate activity (see Figure 3). Finally, the bottom panel depicts galaxies classified based on activity indicated by [O III]/ vs. [O I]/ (see Figure 4). For all three panels, projected cluster-centric distance was normalized to (i.e., ), the radius of a sphere whose density is 200 times the critical density of the universe (cf. Barkhouse et al., 2007).
@@
For a complementary analysis of stellar population age and star formation activity, including Dn(4000) and specific star formation rate metrics, we refer to Paper I (Barkhouse et al., 2025) which explores the UV and IR properties of our galaxy sample. The Dn(4000) metric (Balogh et al., 1999) is a spectral line index constructed from the ratio of the red (measured between 4000 and 4100 Å) to the blue continuum (3850 to 3950 Å). This ratio is sensitive to stellar age and star formation such that a larger value of Dn(4000) implies a greater stellar population age and a lower star formation rate. The star formation rate is obtained from H spectral line measurements and converted to luminosity using galaxy redshift information. The star formation rate is given by SFR , where the H luminosity, L (H), is in erg s-1 (Bell & Kennicutt, 2001; Dhiwar et al., 2023). The specific star formation rate, the star formation rate normalized to galaxy mass, is obtained by dividing the SFR by mass estimated from the extinction- and k-corrected absolute magnitude of each galaxy (Mahajan et al., 2018).
@@ 5 Conclusions
A majority of red spiral galaxies (70.7% to 97.0%, depending on the diagnostic, see Table 1) display AGN or LINER signatures in optical diagnostics, consistent with feedback-driven suppression of star formation.
@@
A significant fraction of blue spirals show high X-ray luminosities, with 40.9% (9 out of 22) having erg s-1 and 22.7% (5 out of 22) exceeding erg s-1, despite being classified as star-forming or composite (star-forming/AGN) in optical diagrams, indicating hidden or optically obscured AGN activity.
@@
The combined use of BPT, WHAN, and extended diagnostics (e.g., Mazzolari et al., 2024, 2025; Backhaus et al., 2025) improves AGN identification by recovering additional AGNs in galaxies where one or more classical BPT lines are weak or undetected. For galaxies with most of the emission-line coverage, the classical BPT diagram identifies AGN/LINER activity in 85% of red spirals and 21% of blue spirals, while extended low-metallicity and low-S/N diagnostics increase the recovered AGN fraction to 70–100% in red spirals and 12–100% in blue spirals, demonstrating that extended diagnostics are more efficient for identifying AGNs in emission-line-weak and low-metallicity systems, whereas the classical BPT diagram remains robust for strong-line galaxies.
@@ Acknowledgments
We thank the reviewer for providing thoughtful comments and suggestions which improved the manuscript.
Appendix A Comparison of X-ray flux measurements
In Figure 1 we show the comparison of X-ray flux measured for three blue spirals and nine red spirals observed by both Chandra and XMM-Newton telescopes. Flux measurements agree to within for 10 of the 12 spirals. For two red spirals, the flux differs by and .